Solve the following system of equations for z and for y System of Equations: Value of Value of y y 9+32 Solve the following system of equations for a and for b: System of Equations: Value of a Value of b 4а + 2b 20 15a +5b 60 Plot the following system of equations on the following graph. System of Equations: p-6-2q

Note: Use the orange line (square symbols) to plot the frst equation, and use the blue line (circle symbols) to plot the second equation. 10 p 6-2q a 5 0 1 2 3 4 5 6 78 9 10

nology kn Ed wsU eHome WSU Information Te Lynda Winona State Univer C) GitHub p 8-2q 6 0 The values of p and q that solve these two equations simultaneously can be seen on the graph as Save & Contir Grade It Now Continue without s
Answer #1

9ナ32 = 65-52 56 =30 Isa +5b = 60 a= 5-0.56 ut valuo ISD 6

10 9 8 7 6 5 4 3 2 1 0

Intersecting point.